Two cyclist, 90 mi apart, start riding toward each other at the same time. One cycles twice as fast as the other. If they meet 2 h later, at what average speed is each cyclist traveling?
Answer by mananth(16949)

Two cyclist, 90 mi apart, start riding toward each other at the same time. One cycles twice as fast as the other. If they meet 2 h later, at what average speed is each cyclist traveling?
..
let the speed of one cyclist be x mph
the other will be 2x mph
..
They both travel for 2 hours before they meet.
..
distance traveled by one + distance traveled by other = Total distance between them when starting.
..
2x+2*2x=90
2x+4x=90
6x=90
x=15 mph the speed of first cyclist.
The speed of second cyclist will be 2x = 2*15=30 mph